A man who is 3 years elder than his wife and 4 times elder than his son.After 3 yrs the age of his son will be 15 so what will be the age of his wife ?

A.42
B.45
C.48
D.60

Answer by Boreal(15235)

x=wife's age
x+3= man's age since he is 3 years older
(x+3)=4* son's age. The son is 12, because in 3 years he will be 15.
x+3=4*12=48
The man is 48.
Now, the wife is 45, but the question is in the future tense. What WILL be the age of his wife? In 3 years, his wife will be 48. The answer is C.